OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I will fan and give a medal!!

OpenStudy (abbycross167):

Please help me to simplify this rational expression - 2xy 4x^2 ----- / ----- 3 3 **Note the / means to divide

OpenStudy (alirayyan):

y/2x

OpenStudy (daniel.ohearn1):

When you divide by a fraction, you can take for example a/b divided by c/d that is equivalent to multiplying by the reciprocal. Therefore a/b / c/d = a/b * d/c = ad/bc

OpenStudy (daniel.ohearn1):

In your case, simply use what you have going.. (2xy/3) / (4x^2 / 3) = (2xy/3)*(3/ 4x^2)

OpenStudy (daniel.ohearn1):

Multiply through and what do you get?

OpenStudy (abbycross167):

@daniel.ohearn1 I'm not so show what the answer to this (2xy/3) / (4x^2 / 3) = (2xy/3)*(3/ 4x^2) would be...